# A Large Language Model Evolutionary Algorithm for Automatically Generating Metaheuristics

### v2

@@ VI-A Performance Analysis in Higher Dimensions
The best-found algorithms using the LLaMEA framework are evaluated against the most relevant state-of-the-art optimizers, namely the previously used baselines CMA-best, CMA-ES and Differential Evolution (DE), CMA-ES and DE are using recommended hyper-parameter settings. The baselines all originate from the IOH analyzer benchmark data set [45]. In Figure 9, the empirical attainment function of the proposed algorithms and baselines is shown for dimension $d\in\{10,20\}$ using results from all BBOB functions, $5$ instances per function and $5$ random seeds ($25$ runs per BBOB function in total). It is interesting to observe that while ERADS performs well in $5$ dimensional problems, it fails to generalize well to higher dimensions, while EnhancedFireworkAlgorithmWithLocalSearch and QuantumDifferentialParticleOptizerWithElitism both have very good performance in higher dimensions for the first $2000$ function evaluations (even better than CMA-best), but around $10000$ function evaluations CMA-best and CMA-ES outperform the other algorithms clearly. When looking at convergence curves per function, we note that this is mainly due to the very hard functions in the Benchmark ($23$ and $24$), See Appendix II for the detailed convergence curves per BBOB function in $5d$, $10d$ and $20d$.
@@
Figure 9: The empirical attainment function (EAF) estimates the percentage of runs that attain an arbitrary target value not later than a given runtime. EAF for the best algorithm per configuration and the baselines; CMA-ES, CMA-best and DE, averaged over all 24 BBOB functions in (from top to bottom) $10d$ and $20d$ respectively.
@@ VII Conclusions and Outlook
This paper introduced LLaMEA, a novel framework leveraging Large Language Models (LLMs) for the automatic generation and optimization of metaheuristic algorithms. Our approach uniquely automates the evolution of algorithm design, enabling efficient exploration and optimization within a computationally feasible framework. Our findings demonstrate that LLaMEA can effectively generate high-performing algorithms that rival and sometimes surpass existing state-of-the-art techniques.
The LLaMEA framework proved capable of generating and evolving algorithms that perform comparably to traditional state-of-the-art metaheuristics, demonstrating the potential of LLMs to understand and innovate within the algorithmic design space effectively. Algorithms evolved through our framework successfully compete against established metaheuristics, highlighting the practical applicability of using LLMs for automated construction of optimization heuristics. Since LLMs have been trained on an exceedingly large code base, including the currently available metaheuristics, we can explain their observed strong performance by interpreting them as a universal modular framework for algorithm construction (see also Section I) that have access to a huge number of ”modules” that can be combined to form new algorithms.
@@
  • Future work could explore the expansion of the LLaMEA framework to support a broader range of evolutionary strategies. While the proposed framework focuses on an (1 $\stackrel{{\scriptstyle+}}{{,}}\,$1)-EA approach, where we have one parent and generate one solution at a time, it is possible to generalize the framework to a ($\mu\,\stackrel{{\scriptstyle+}}{{,}}\,\lambda$)-EA, keeping a larger population and generating multiple individuals. This would translate to generating multiple mutations with LLMs by leveraging multiple random seeds in the generation process.
@@
  • Applying the LLaMEA methodology to other domains such as automated machine learning (AutoML) and complex system design could further demonstrate the versatility and impact of this approach.
@@
Reproducibility statement. We provide an open-source documented implementation of our package at [31], including install and how-to-use guide.
Intermediate results, generated algorithms and BBOB evaluation results are available as well in subdirectories of the repository.

### v3

@@ VI-A Performance Analysis in Higher Dimensions
The best-found algorithms using the LLaMEA framework are evaluated against the most relevant state-of-the-art optimizers, namely the previously used baselines CMA-best, CMA-ES and Differential Evolution (DE). CMA-ES and DE are using recommended hyperparameter settings. The baselines all originate from the IOHanalyzer benchmark data set [54]. In Figure 9, the empirical attainment function of the proposed algorithms and baselines are shown for dimension $d\in\{10,20\}$ using results from all BBOB functions, $5$ instances per function and $5$ random seeds ($25$ runs per BBOB function in total). It is interesting to observe that while ERADS performs well in $5$ dimensional problems, it fails to generalize well to higher dimensions, while EnhancedFireworkAlgorithmWithLocalSearch and QuantumDifferentialParticleOptizerWithElitism both have very good performance in higher dimensions for the first $2000$ function evaluations (even better than CMA-best), but around 10 0001000010\,00010 000 function evaluations CMA-ES outperforms the other algorithms clearly. It is also interesting to observe that the optimized CMA-best performs less well than a CMA-ES with default hyperparameters in $10d$ and $20d$. Just like the LLaMEA proposed algorithms, CMA-ES best was optimized on $5d$ and with a budget of 10 0001000010\,00010 000 evaluations. When looking at convergence curves per function, we note that CMA-ES outperforms mainly on the very hard functions in the benchmark ($f_{23}$ and $f_{24}$). Detailed convergence curves per BBOB function in $5d$, $10d$ and $20d$ are available in the supplemental material [39].
We would like to emphasize that the goal of our research is to show that, for a specific setting (here: BBOB in $5d$), an LLM can generate a superior algorithm automatically. This algorithm is not expected to scale to high-dimensional problems, as it was generated for the specific case of $5d$ problems. For this reason, we do not go beyond $d=20$ for further evaluation of the generated algorithms.
@@
Figure 9: The empirical attainment function (EAF) estimates the percentage of runs that attain an arbitrary target value not later than a given runtime. EAF for the best algorithm per configuration and the baselines; CMA-ES, CMA-best and DE, averaged over all 24 BBOB functions in $10d$ (left plot) and $20d$ (right plot), respectively.
@@ VII Conclusions and Outlook
This paper introduced LLaMEA, a novel framework leveraging Large Language Models (LLMs) for the automatic generation and optimization of metaheuristic algorithms. Our approach automates the evolution of algorithm design, enabling efficient exploration and optimization within a computationally feasible framework. Our findings demonstrate that LLaMEA can effectively generate high-performing algorithms that rival and sometimes surpass existing state-of-the-art techniques.
The LLaMEA framework proved capable of generating and evolving algorithms that perform comparably to traditional state-of-the-art metaheuristics, demonstrating the potential of LLMs to understand and innovate within the algorithmic design space effectively. Algorithms evolved through our framework successfully compete against established metaheuristics, highlighting the practical applicability of using LLMs for automated construction of optimization heuristics. Since LLMs have been trained on an exceedingly large code base, including the currently available metaheuristics, we can explain their observed strong performance by interpreting them as a universal modular framework for algorithm construction (see also Section I) that have access to a huge number of “modules” that can be combined to form new algorithms. In addition, we observed that the LLM is able to both fine-tune algorithm parameters, as well as introduce new logic such as different mutation and crossover strategies in the generated algorithms.
Considering how much work goes into the manual design of ”nature-inspired heuristics”, which are often not novel [55, 56, 57, 58] and often cannot compete with random search [10], we argue that an automated design approach for specific application domains will likely be the method of choice from now on.
@@
  • Future work could explore the expansion of the LLaMEA framework to support a broader range of evolutionary strategies. While the proposed framework focuses on an (1 $\stackrel{{\scriptstyle+}}{{,}}\,$1)-EA approach, where we have one parent and generate one solution at a time, it is possible to generalize the framework to a ($\mu\,\stackrel{{\scriptstyle+}}{{,}}\,\lambda$)-EA (i.e., an algorithm with parent population size $\mu$ and offspring population size $\lambda\gg\mu$), as in population-based evolutionary algorithms [2], keeping a larger population and generating multiple individuals. This would translate into generating multiple mutations and recombinations with LLMs by leveraging multiple random seeds and different temperature values in the generation process.
@@
  • Applying the LLaMEA methodology to other algorithm classes for continuous optimization problems, such as Bayesian and more general surrogate-assisted algorithms, could further demonstrate the versatility of LLaMEA.
@@
Reproducibility statement. We provide an open-source documented implementation of our package at [39]. Intermediate results, generated algorithms and BBOB evaluation results are available as well in subdirectories of the repository.
Supplementary material
TABLE II: The 24 BBOB noiseless functions grouped in five function categories, see [30] for details.
Group Function ID Function Name
1 f1 Sphere Function
f2 Separable Ellipsoidal Function
f3 Rastrigin Function
f4 Büche-Rastrigin Function
f5 Linear Slope
2 f6 Attractive Sector Function
f7 Step Ellipsoidal Function
f8 Rosenbrock Function, original
f9 Rosenbrock Function, rotated
3 f10 Ellipsoidal Function
f11 Discus Function
f12 Bent Cigar Function
f13 Sharp Ridge Function
f14 Different Powers Function
4 f15 Rastrigin Function
f16 Weierstrass Function
f17 Schaffer’s F7 Function
f18 Schaffer’s F7 Function, moderately ill-conditioned
f19 Composite Griewank-Rosenbrock Function F8F2
5 f20 Schwefel Function
f21 Gallagher’s Gaussian 101-me Peaks Function
f22 Gallagher’s Gaussian 21-hi Peaks Function
f23 Katsuura Function
f24 Lunacek bi-Rastrigin Function
Figure 10: Median best-so-far function value ($y$-axis) over the 2 $2\,000\cdot d$ function evaluations ($x$-axis) per BBOB function for the best algorithm per configuration and the baselines; CMA-ES, DE and BIPOP-CMA-ES (for $d=5$).
Figure 11: Median best-so-far function value ($y$-axis) over the 2 $2\,000\cdot d$ function evaluations ($x$-axis) per BBOB function for the best algorithm per configuration and the baselines; CMA-ES, DE and BIPOP-CMA-ES (for $d=10$).
Figure 12: Median best-so-far function value ($y$-axis) over the 2 $2\,000\cdot d$ function evaluations ($x$-axis) per BBOB function for the best algorithm per configuration and the baselines; CMA-ES, DE and BIPOP-CMA-ES (for $d=20$).
Figure 13: A code diff example snippet to show how the LLM mutates the parent into an offspring. This diff shows a variation of hyper-parameters (tuning) and the arguments in comments the LLM uses to change them, including a new mutation strategy. This change occurred at the beginning of the evolution process (early iteration).
Figure 14: A code diff example snippet to show how the LLM mutates the parent into an offspring. This diff shows how the LLM fine-tunes some of the hyper-parameters, this change occurred at the end of the evolution process, showing a small mutation (the remaining 30 lines of code, line 34 and further, were unchanged).
Figure 15: A code diff example snippet to show how the LLM mutates the parent into an offspring. This diff varies the crossover operator of the algorithm to a new approach.